# Low temperature ferroelectric state in strontium titanate microcrystals using in situ multi-reflection Bragg coherent X-ray diffraction imaging

### v1

@@ Abstract
Strontium titanate is a classic quantum paraelectric oxide material that has been widely studied in bulk and thin films. It exhibits a well-known cubic-to-tetragonal antiferrodistortive phase transition at 105 K, characterized by the rotation of oxygen octahedra. A possible second phase transition at lower temperature is suppressed by quantum fluctuations, preventing the onset of ferroelectric order. However, recent studies have shown that ferroelectric order can be established at low temperatures by inducing strain and other means. Here, we used in situ multi-reflection Bragg coherent X-ray diffraction imaging to measure the strain and rotation tensors for two strontium titanate microcrystals at low temperature. We observe strains induced by dislocations and inclusion-like impurities in the microcrystals. Based on radial magnitude plots, these strains increase in magnitude and spread as the temperature decreases. Pearson’s correlation heatmaps show a structural transition at 50 K, which we associate with the formation of a low-temperature ferroelectric phase in the presence of strain. We do not observe any change in local strains associated with the tetragonal phase transition at 105 K.
@@ I Introduction
At low temperatures, STO undergoes notable changes in its structure and properties. At 105 K, it experiences a well-known cubic-to-tetragonal antiferrodistortive (AFD) phase transition [4] characterized by the rotation of the TiO6 octahedra [5, 6, 7], resulting in a of 1.0009 for the tetragonal unit cell at 10 K [8]. While there is no abrupt phase transition identified, there are several reports of anomalous behavior below 50K. The thermal expansion coefficient starts changing around that temperature [9]. Below 50 K, it exhibits high dielectric constants that deviate from the classical Curie-Weiss law [10]. STO is a classic quantum paraelectric material, meaning it remains paraelectric at low temperatures. The polar modes in such materials soften as the temperature decreases [7], approaching a phase transition but never fully becoming ferroelectric due to suppression from quantum fluctuations [11]. These quantum fluctuations stabilize the paraelectric phase, preventing the onset of long-range ferroelectric order.
@@
However, the presence of this quantum suppression supports the description of STO as an incipient ferroelectric. In the ferroelectric phase, the Ti ions shift from their central positions in the oxygen octahedra, breaking the centrosymmetry and creates a dipole moment within the crystal. However, many have suggested that this movement is subdued by the AFD octahedra rotations [12, 13]. Nonetheless, ferroelectricity can be introduced by small perturbations such as strain [14], isotope substitution [15], Ca2+ substitution [16], electric fields [6, 17], and laser excitation [18, 19]. Furthermore, the low-temperature phase is positioned close to a quantum critical point, which is related to superconductivity [20]. Understanding the physics behind these phenomena is crucial for the development of next-generation electronic devices.
@@
Many low-temperature studies have been performed on bulk single crystal STO [10, 9, 11, 6, 4, 7, 21, 15, 17, 18, 20] or thin films [22, 23, 14]. Investigating the properties of functional materials has recently been focused on the sub-micron level, where surface energy and sample morphology can also dictate the material’s properties, emphasizing the need for further exploration. STO microcrystals can be fabricated using solid-state synthesis [24] and the sol-gel method [25], and are widely used in catalysis [1]. We synthesized STO microcrystals (Fig. 1) using hydrothermal synthesis [26], a technique that leverages the solubility differences of metal precursors at elevated temperatures and pressures, leading to the formation of highly crystalline samples that vary in size based on incubation duration. Of particular interest is the role of strain in dictating the properties of STO microcrystals. For thin film samples, the ferroelectric behavior of STO can be influenced by strain caused by dopants [27] or epitaxy [14]. However, this behavior has not been previously measured in free-standing microcrystals.
@@
To probe the strain in microcrystals, we use a non-destructive technique called Bragg coherent X-ray diffraction imaging (BCDI) [28, 29]. BCDI allows for three-dimensional (3D) imaging of isolated micro- and nanocrystalline materials with a spatial resolution of up to ten nanometres. It can provide the 3D lattice strain field, projected along the scattering vector, , of a particular crystal reflection, with a resolution on the order of [30, 31]. BCDI has been recently used for many in situ experiments including ZrO grain growth [32], Pt catalysis [33], Li- and Mn-rich cathode degradation [34], and barium titanate phase transformation [35].
@@
BCDI involves fully illuminating a crystalline sample inside the coherent volume of an X-ray beam, which is typically no larger than in any dimension at third- and fourth-generation synchrotron sources [29]. Once the Bragg condition is met for a specific reflection, the diffraction pattern is collected on a pixelated area detector positioned perpendicular to the outgoing wave vector in the Fraunhofer diffraction regime. By rotating the sample about a rocking axis, a 3D coherent X-ray diffraction pattern (CXDP) is collected as the detector moves through the Ewald sphere. If the CXDP is oversampled by at least twice the Nyquist frequency [36], at least 2 pixels per fringe along one dimension, iterative phase retrieval algorithms that apply constraints in real and reciprocal space until convergence is satisfied, can be used to recover the phase [37, 38]. The amplitude and phase in reciprocal space are related to the real-space object via an inverse Fourier transform [39, 28], followed by a real space transformation from detector-conjugated space to orthogonal lab or sample space [40]. The resulting amplitude, , where is the position vector, is proportional to the effective electron density of the crystalline volume associated with the particular crystal reflection. The phase, , corresponds to the projection of the lattice displacement field, , onto the Bragg vector:
@@
If at least three linearly independent reflections are collected from a single microcrystal, known as multi-reflection BCDI (MBCDI), one can compute the 3D lattice strain and rotation tensors with respect to an arbitrary reference (section II.4). Here, we set the center of mass of the reconstructed crystal for each reflection as the zero reference. MBCDI requires knowledge of the crystal orientation, which is typically attained via estimates based on a priori crystal geometry [41], synchrotron micro-beam Laue diffraction [42, 43], indexing pole figures corresponding to a known Bragg peak [44], and/or laboratory-based electron backscatter diffraction (EBSD) [45]. Obtaining multiple reflections for a single crystal can be challenging and labor-intensive, but has been valuable for studying defects and dislocations [31, 46], domain structures [47] and enabling simultaneous MBCDI phase retrieval procedures to increase reconstruction fidelity [48, 49, 50, 51].
@@
Here, we study two STO microcrystals, referred to as crystal A and crystal B, at low temperatures. We analyze different structural features in each microcrystal. The first crystal has two dislocations present, which contribute to strain fields that evolve with temperature. The second crystal contains reproducible regions of low electron density, which we refer to as voids, suggesting there is nonuniformity present. Despite these differences, there are temperature trends that remain consistent for both crystals. Although this has recently been accomplished at high temperatures [52], we use in situ MBCDI to determine the residual strain and rotation tensors at cryogenic temperatures, allowing us to capture subtle structural changes. We use heatmaps to examine changes in structure in real and reciprocal space, enabling the detection of different structural modes of STO. Furthermore, the 3D tensor information is analyzed as radial averages, allowing for qualitative monitoring of the strain patterns.
@@ II.1 Hydrothermal synthesis
The synthesis of the STO microcrystals with a cubic shape was inspired by the work of Hao et al. [25]. It first involved the preparation of TiCl4, SrCl2, and LiOH (3M) solutions. Strontium chloride (hexahydrate, 99%), lithium hydroxide (monohydrate, 98%), and titanium (IV) chloride (99.9%) were sourced from Sigma-Aldrich. The TiCl4 solution involved adding 0.26 mL TiCl4 (cooled overnight in a fridge) to a mixture of 10 mL MeOH and 25 mL H2O. The SrCl2 solution was prepared by adding 0.24 g SrCl2 to 10 mL H2O. 3.77 g LiOH was added into 30 mL H2O to form LiOH 3M. Next, we added the TiCl4 and SrCl2 mixture into 3M LiOH to create a white suspension in a Teflon vessel, which was stirred for 30 minutes at room temperature. The sealed Teflon vessel was then positioned in a stainless steel autoclave and treated at 200°C for 48 hours. After, the autoclave was allowed to cool down to room temperature, and the sub-micron crystals were then centrifuged and washed several times with acetone to obtain the powders. The STO particles were then immersed in a 0.035 wt.% solution of poly(ethyleneimine) in distilled water. This solution was drop-cast onto a silicon substrate, which was then calcined in a furnace at 550°C for 4 hours to complete the sampling process for BCDI measurement. Crystals A and B were from this batch of crystals.
@@ II.2 BCDI data collection
We accomplished this for two crystals. For crystal A, we measured the , , , , and Bragg peaks at 10 K, 20 K, 30 K, 40 K, 50 K, 100 K, 150 K, 200 K, 250 K, and 300 K. However, the reconstruction for the Bragg peak did not converge as well and was therefore excluded from the analysis, except for the determination of the Burgers vector (Table 1). For crystal B, the and Bragg peaks were measured at 180 K, 160 K, 140 K, 100 K, 80 K, 50 K, 40 K, 30 K, 20 K and 10 K, and the Bragg peak was measured at 180 K, 160 K, 140 K, 100 K, 80 K, 50 K. The temperature ramp rate was 5 K/min, and once a specific temperature was reached, 5 minutes elapsed before starting the BCDI measurements. The sample temperature was measured using a calibrated diode attached to the copper stub. The sample was kept in the beam as the cryostat thermally contracted by correcting the height using a previously measured calibration curve.
Coherent X-ray diffraction patterns (CXDPs) were collected on a 512 × 512 pixel Quad-Merlin detector with a pixel size of , positioned at 1.31 m from the sample to ensure oversampling. CXDPs were recorded by rotating the crystal through an angular range of 0.5 - 0.6° and recording an image every 0.005° with a 1 s exposure time. To increase the signal-to-noise ratio, five repeated scans for each of the reflections at every temperature were measured.
@@ II.3 Data processing
The CXDP was cropped to a size of pixels in the detector plane to avoid detector module gaps during data processing. Each set of five repeated scans was aligned using a 3D version of a 2D subpixel translation procedure proposed by Guizar-Sicairos et al. [53]. After alignment, the Pearson correlation of the set of repeated CXDPs was greater than 0.99, demonstrating that the crystal and environmental setup remained remarkably stable in the cryostat for most of the measurements. If the Pearson correlation of a repeated scan was below 0.99, it was excluded. The scans were summed, and the background was determined as the mean value of a fitted histogram of the summed data. Before fitting the histogram, values below 1 and those greater than 95% of the maximum intensity were removed. The background of the five summed scans was subtracted from the CXDP and also used as the minimum data threshold to recover the complex electron density in the phase retrieval algorithm. Further details can be found in Appendix A.
@@ III.1 Crystal A with dislocations
The tensor components and average morphology for crystal A are shown in Fig. 2. This crystal is 400 nm in length and exhibited lattice strain patterns in the and planes, indicating the presence of dislocations [31]. This observation was further confirmed by the phase vortices in each measured reflection, shown in Appendix B.
@@ III.1.1 Crystal A dislocation identification
We created a “dislocation mask”, shown as a dark region surrounding each dislocation in Fig. 3. This region was created by masking a 90 nm radius along each average dislocation line and covered roughly 25% of the average crystal morphology volume. Its purpose was to exclude most of the lattice strain contributions from the dislocation core, as discussed in section III.1.3.
@@ III.1.2 Crystal A Pearson correlation coefficient
where and were the temperatures being compared, and were the values for a single voxel, and and were the means of each array. Prior to computing Pearson’s coefficient for the CXDPs, we processed them based on the procedure listed in section II.3. Afterward, the CXDPs were binarized by setting values greater than zero to one, thereby ignoring the intensity values for Pearson correlation comparison. Fig. 4 shows the trends associated with each reflection and temperature for crystal A.
@@ III.1.3 Crystal A radial strain and rotation tensor magnitudes
Using the extreme ends of the temperature measurements, 10 K and 300 K, we compared the differences in the tensor components at these two temperature states. To quantify the strain distribution within the crystal at different temperatures, the magnitude of the tensor component was averaged over successively larger shells from the center of the crystal [55, 62]. For each radial position, the average magnitude of the tensor components, along with the standard deviation of the magnitudes, is plotted in Fig. 6(a). In Fig. 6(b) and 6(c), this distribution was subdivided into regions of strain located close to the dislocations and far away, by use of the dislocation mask shown in Fig. 3.
@@
The magnitude is slightly higher at 10 K compared to 300 K, especially close to the center of the crystal where no dislocations were present. To further analyze this, we computed the relative volume change, or volumetric strain, , defined as .
@@ III.2 Crystal B with voids
To confirm the results in Section III.1, a second crystal from the same synthesis batch was measured in a separate experiment at different temperatures and for different reflections. The tensor components and average morphology for crystal B are shown in Fig. 8. The individual reflections used for this computation are discussed in Appendix B. This crystal is 600 nm in length and, unlike crystal A, showed no signs of dislocations. Rather, crystal B contained spherical-shaped regions where the effective electron density falls well below that of the rest of the crystal, which we consider to be “voids”. These voids, highlighted as colored circles in Fig. 8 in the plane, appeared in slightly different locations for different reflections, as indicated by the colors.
Fig. 9 shows a translucent overlap of the normalized amplitude for the three measured reflections. Despite strong overlap along the surfaces of each reconstruction, there is poor overlap for the voids at each temperature. However, the voids for each reflection have relatively consistent positions across different temperatures, though their sizes are slightly different. The size differences were due to minor variations in the reconstructed amplitude surrounding the voids. The lack of electron density information in the reconstructions implies that the void regions were either not diffracting under the given Bragg conditions or the information was not captured by the detector.

### v4

# Low temperature state in strontium titanate microcrystals using in situ multi-reflection Bragg coherent x-ray diffraction imaging

@@ Abstract
Strontium titanate is a classic quantum paraelectric oxide material that has been widely studied in bulk and thin films. It exhibits a well-known cubic-to-tetragonal antiferrodistortive phase transition at 105 K, characterized by the rotation of oxygen octahedra. A possible second phase transition at lower temperature is suppressed by quantum fluctuations, preventing the onset of ferroelectric order. However, recent studies have shown that ferroelectric order can be established at low temperatures by inducing strain and other means. Here, we used in situ multi-reflection Bragg coherent x-ray diffraction imaging to measure the strain and rotation tensors for two strontium titanate microcrystals at low temperature. We observe strains induced by dislocations and inclusion-like impurities in the microcrystals. Based on radial magnitude plots, these strains increase in magnitude and spread as the temperature decreases. Pearson’s correlation heat maps show a structural transition at 50 K, which could possibly be the formation of a low-temperature ferroelectric phase in the presence of strain. We do not observe any change in local strains associated with the tetragonal phase transition at 105 K.
@@ I Introduction
At low temperatures, STO undergoes notable changes in its structure and properties. At 105 K, it experiences a well-known cubic-to-tetragonal antiferrodistortive (AFD) phase transition [4] characterized by the rotation of the TiO6 octahedra [5, 6, 7], resulting in a of 1.0009 for the tetragonal unit cell at 10 K [8]. While there is no abrupt phase transition identified, there are several reports of interesting behavior below 50K. The thermal expansion coefficient starts changing around that temperature [9]. Below 50 K, it exhibits high dielectric constants that deviate from the classical Curie-Weiss law [10]. STO is a classic quantum paraelectric material, meaning it remains paraelectric at low temperatures. The polar modes in such materials soften as the temperature decreases [7], approaching a phase transition but never fully becoming ferroelectric due to suppression from quantum fluctuations [11]. These quantum fluctuations stabilize the paraelectric phase, preventing the onset of long-range ferroelectric order.
@@
However, the presence of this quantum suppression supports the description of STO as an incipient ferroelectric. In the ferroelectric phase, the Ti ions shift from their central positions in the oxygen octahedra, breaking the centrosymmetry and creating a dipole moment within the crystal. However, many have suggested that this movement is subdued by the AFD octahedra rotations [12, 13]. Nonetheless, ferroelectricity can be introduced by small perturbations such as strain [14], isotope substitution [15], Ca2+ substitution [16], electric fields [6, 17], and laser excitation [18, 19]. Furthermore, the low-temperature phase is positioned close to a quantum critical point, which is related to superconductivity [20]. Understanding the physics behind these phenomena is crucial for the development of next-generation electronic devices.
@@
Many low-temperature studies have been performed on bulk single crystal STO [10, 9, 11, 6, 4, 7, 21, 15, 17, 18, 20] or thin films [22, 23, 14]. Investigating the properties of functional materials has recently been focused on the sub-micron level, where surface energy and sample morphology can also dictate the material’s properties, emphasizing the need for further exploration. STO microcrystals can be fabricated using solid-state synthesis [24] and the sol-gel method [25], and are widely used in catalysis [1]. We synthesized STO microcrystals (Fig. 1) using hydrothermal synthesis [26], a technique that leverages the solubility differences of metal precursors at elevated temperatures and pressures, leading to the formation of highly crystalline samples that vary in size based on incubation duration and temperature. Of particular interest is the role of strain in dictating the properties of STO microcrystals. For thin film samples, the ferroelectric behavior of STO can be influenced by strain caused by dopants [27] or epitaxy [14]. However, this behavior has not been previously measured in free-standing microcrystals.
@@
To probe the strain in microcrystals, we use a non-destructive technique called Bragg coherent x-ray diffraction imaging (BCDI) [28, 29]. BCDI allows for three-dimensional (3D) imaging of isolated micro- and nanocrystalline materials with a spatial resolution of up to 10 nm. It can provide the 3D lattice strain field, projected along the scattering vector, , of a particular crystal reflection, with a resolution on the order of [30, 31]. BCDI has been recently used for many in situ experiments including ZrO grain growth [32], Pt catalysis [33], Li- and Mn-rich cathode degradation [34], and barium titanate phase transformation [35].
@@
BCDI involves fully illuminating a crystalline sample inside the coherent volume of an x-ray beam, which is typically no larger than in any dimension at third- and fourth-generation synchrotron sources [29]. Once the Bragg condition is met for a specific reflection, the diffraction pattern is collected on a pixelated area detector positioned perpendicular to the outgoing wave vector in the Fraunhofer diffraction regime. By rotating the sample about a rocking axis, a 3D coherent x-ray diffraction pattern (CXDP) is collected as the detector moves through the Ewald sphere. If the CXDP is oversampled by at least twice the Nyquist frequency [36], at least 2 pixels per fringe along one dimension, iterative phase retrieval algorithms that apply constraints in real and reciprocal space until convergence is satisfied, can be used to recover the phase [37, 38]. The amplitude and phase in reciprocal space are related to the real-space object via an inverse Fourier transform [39, 28], followed by a real space transformation from detector-conjugated space to orthogonal laboratory or sample space [40]. The resulting amplitude, , where is the position vector, is proportional to the effective electron density of the crystalline volume associated with the particular crystal reflection. The phase, , corresponds to the projection of the lattice displacement field, , onto the Bragg vector:
@@
If at least three linearly independent reflections are collected from a single microcrystal, known as multi-reflection BCDI (MBCDI), one can compute the 3D lattice strain and rotation tensors with respect to an arbitrary reference (Sec. II.4). Here, we set the center of mass of the reconstructed crystal for each reflection as the zero reference. MBCDI requires knowledge of the crystal orientation, which is typically attained via estimates based on a priori crystal geometry [41], synchrotron micro-beam Laue diffraction [42, 43], indexing pole figures corresponding to a known Bragg peak [44], and/or laboratory-based electron backscatter diffraction (EBSD) [45]. Obtaining multiple reflections for a single crystal can be challenging and labor-intensive, but has been valuable for studying defects and dislocations [31, 46], domain structures [47] and enabling simultaneous MBCDI phase retrieval procedures to increase reconstruction fidelity [48, 49, 50, 51].
@@
Here we explore how defects affect STO microcrystals at low temperature, and how it might differ compared to bulk () or thin film STO. We note that some fundamental studies on the low temperature phase in bulk STO have possibly been influenced by sample damage near the surface of the crystal, which contain defects [52, 7, 53, 54]. With MBCDI, we isolate how these defects affect the low temperature behavior of STO. We study two STO microcrystals, referred to as crystal A and crystal B, at low temperatures. We analyze different structural features in each microcrystal. The first crystal has two dislocations present, which contribute to strain fields that evolve with temperature. The second crystal contains reproducible regions of low electron density, which we refer to as voids, suggesting there is nonuniformity present. Despite these differences, there are temperature trends that remain consistent for both crystals. Although this has recently been accomplished at high temperatures [55], we use in situ MBCDI to determine the residual strain and rotation tensors at cryogenic temperatures, allowing us to capture subtle strain changes () surrounding defects. We use heat maps to examine changes in structure in real and reciprocal space, enabling the detection of different structural modes of STO. Furthermore, the 3D tensor information is analyzed as radial averages, allowing for qualitative monitoring of the strain patterns.
@@ II.1 Hydrothermal synthesis
The synthesis of the STO microcrystals with a cubic shape was inspired by the work of Hao et al. [25]. It first involved the preparation of TiCl4, SrCl2, and LiOH (3M) solutions. Strontium chloride (hexahydrate, 99%), lithium hydroxide (monohydrate, 98%), and titanium (IV) chloride (99.9%) were sourced from Sigma-Aldrich. The TiCl4 solution involved adding 0.26 mL TiCl4 (cooled overnight in a fridge) to a mixture of 10 mL MeOH and 25 mL H2O. The SrCl2 solution was prepared by adding 0.64 g SrCl2 to 10 mL H2O. 3.77 g LiOH was added into 30 mL H2O to form 3M LiOH. Next, we added the TiCl4 and SrCl2 mixture into 3M LiOH to create a white suspension in a Teflon vessel, which was stirred for 30 minutes at room temperature. The sealed Teflon vessel was then positioned in a stainless steel autoclave and treated at 200 °C for 48 h. After, the autoclave was allowed to cool down to room temperature, and the sub-micron crystals were then centrifuged and washed several times with acetone to obtain the powders. The STO particles were then immersed in a 0.035 wt.% solution of poly(ethyleneimine) in distilled water. This solution was drop-cast onto a silicon substrate, which was then calcined in a furnace at 550 °C for 4 h to complete the sampling process for BCDI measurement. Crystals A and B were from this batch of crystals.
@@ II.2 BCDI data collection
We accomplished this for two crystals, which were selected based on having strong Bragg peak signals. For crystal A, we measured the , , , , and Bragg peaks at 10, 20, 30, 40, 50, 100, 150, 200, 250, and 300 K. However, the reconstruction for the Bragg peak did not converge as well and was therefore excluded from the analysis, except for the determination of the Burgers vector (Table 1). For crystal B, the and Bragg peaks were measured at 180, 160, 140, 100, 80, 50, 40, 30, 20 and 10 K, and the Bragg peak was measured at 180, 160, 140, 100, 80, and 50 K. The temperature ramp rate was 5 K/min, and once a specific temperature was reached, 5 minutes elapsed before starting the BCDI measurements. The sample temperature was measured using a calibrated diode attached to the copper stub. The sample was kept in the beam as the cryostat thermally contracted by correcting the height using a previously measured calibration curve.
Coherent x-ray diffraction patterns (CXDPs) were collected on a 512 × 512 pixel Quad-Merlin detector with a pixel size of , positioned at 1.31 m from the sample to ensure oversampling. CXDPs were recorded by rotating the crystal through an angular range of 0.5° - 0.6° and recording an image every 0.005° with a 1 s exposure time. To increase the signal-to-noise ratio, five repeated scans for each of the reflections at every temperature were measured.
@@ II.3 Data processing
The CXDP was cropped to a size of pixels in the detector plane to avoid detector module gaps during data processing. Each set of five repeated scans was aligned using a 3D version of a 2D subpixel translation procedure proposed by Guizar-Sicairos et al. [56]. After alignment, the Pearson correlation of the set of repeated CXDPs was greater than 0.99, demonstrating that the crystal and environmental setup remained remarkably stable in the cryostat for most of the measurements. The scans were summed, and the background was determined as the mean value of a fitted histogram of the summed data. Before fitting the histogram, values below 1 and those greater than 95% of the maximum intensity were removed. The background of the five summed scans was subtracted from the CXDP and also used as the minimum data threshold to recover the complex electron density in the phase retrieval algorithm. Further details can be found in Appendix A.
@@ III.1 Crystal A with dislocations
The tensor components and average morphology for crystal A are shown in Fig. 2. This crystal is 400 nm in length and exhibited lattice strain patterns in the y-z and x-z planes, indicating the presence of dislocations [31]. This observation was further confirmed by the phase vortices in each measured reflection, shown in Appendix B.
@@ III.1.1 Crystal A dislocation identification
We created a “dislocation mask”, shown as a dark region surrounding each dislocation in Fig. 3. This region was created by masking a 90 nm radius along each average dislocation line and covered roughly 25% of the average crystal morphology volume. Its purpose was to exclude most of the lattice strain contributions from the dislocation core, as discussed in Sec. III.1.3.
@@ III.1.2 Crystal A Pearson correlation coefficient
where and were the temperatures being compared, and were the values for a single voxel, and and were the means of each array. Prior to computing Pearson’s coefficient for the CXDPs, we processed them based on the procedure listed in Sec. II.3. Afterward, the CXDPs were binarized by setting values greater than zero to one, thereby ignoring the intensity values for Pearson correlation comparison. Fig. 4 shows the trends associated with each reflection and temperature for crystal A.
@@ III.1.3 Crystal A radial strain and rotation tensor magnitudes
Using the extreme ends of the temperature measurements, 10 and 300 K, we compared the differences in the tensor components at these two temperature states. To quantify the strain distribution within the crystal at different temperatures, the magnitude of the tensor component was averaged over successively larger shells from the center of the crystal [58, 66]. For each radial position, the average magnitude of the tensor components, along with the standard deviation of the magnitudes, is plotted in Fig. 6(a). In Figs. 6(b) and 6(c), this distribution was subdivided into regions of strain located close to the dislocations and far away, by use of the dislocation mask shown in Fig. 3.
@@
The magnitude is slightly higher at 10 K compared to 300 K, especially close to the center of the crystal where no dislocations are present. To further analyze this, we computed the relative volume change, or volumetric strain, , defined as .
@@ III.2 Crystal B with voids
To confirm the results in Sec. III.1, a second crystal from the same synthesis batch was measured in a separate experiment at different temperatures and for different reflections. The tensor components and average morphology for crystal B are shown in Fig. 8. The individual reflections used for this computation are discussed in Appendix B. This crystal is 600 nm in length and, unlike crystal A, showed no signs of dislocations. Rather, crystal B contained spherical-shaped regions where the effective electron density falls well below that of the rest of the crystal, which we consider to be “voids”. These voids, highlighted as colored circles in Fig. 8 in the y-z plane, appeared in slightly different locations for different reflections, as indicated by the colors.
Fig. 9 shows a translucent overlap of the normalized amplitude for the three measured reflections. Despite strong overlap along the surfaces of each reconstruction, there is poor overlap for the voids at each temperature. However, the voids for each reflection have relatively consistent positions across different temperatures, though their sizes are slightly different. The size differences were due to minor variations in the reconstructed amplitude surrounding the voids. The voids only represent the amplitude region that falls below the threshold of 0.20, but the volume of lower amplitude likely is larger than the void shown in each reflection and can overlap with the other volumes of lower amplitude in the other reflections (see Fig. 18 in Appendix B). The lack of electron density information in the reconstructions implies that the void regions were either not diffracting under the given Bragg conditions or the information was not captured by the detector. Since the voids appear uniquely in certain reflections, one possibility is that the center of crystal B is composed of multiple smaller, crystalline impurities, forming the seed during hydrothermal synthesis. These smaller impurities would diffract under certain Bragg conditions, similar to twin domains within a crystal [67].